### v2

Abstract
Adapting pre-trained deep learning models to new and unknown environments remains a major challenge in underwater acoustic localization. We show that although the performance of pre-trained models suffers from mismatch between the training and test data, they generally exhibit a higher uncertainty in environments where there is more mismatch. Additionally, in the presence of environmental mismatch, spurious peaks can appear in the output of classification-based localization approaches, which inspires us to define and use a method to quantify the “implied uncertainty” based on the number of model output peaks. Leveraging this notion of implied uncertainty, we partition the test samples into sets with more certain and less certain samples, and implement a method to adapt the model to new environments by using the certain samples to improve the labeling for uncertain samples. Thus, using this efficient method for model uncertainty quantification, we showcase an innovative approach to adapt a pre-trained model to unseen underwater environments at test time. This eliminates the need for labeled data from the target environment or the original training data. This adaptation is enhanced by integrating an independent estimate based on the received signal energy. We validate the approach extensively using real experimental data, as well as synthetic data consisting of model-generated signals with real ocean noise. The results demonstrate significant improvements in model prediction accuracy, underscoring the potential of the method to enhance underwater acoustic localization in diverse, noisy, and unknown environments.
I Introduction
Underwater acoustic (UWA) source localization refers to determining all or some of the coordinates of an underwater acoustic source using the corresponding recorded signals [1, 2, 3, 4, 5, 6]. Source ranging refers to estimating the distance of a source and can be achieved using conventional methods such as matched field processing (MFP) [1, 2] or deep learning (DL)-based methods [3, 7, 8]. DL-based UWA localization algorithms [3, 9, 7, 10] tend to generalize poorly to mismatched environments[7, 11, 12]. To overcome small mismatches between the testing and training sets, domain adaptation (DA) methods have been proposed that match some of the statistics of the testing data to those of the training data[13], which, at inference time, requires access to the training data acquired with known or controlled sources. However, for low-power or low storage-capacity underwater devices working in a decentralized manner [14, 15, 16], training data will not always be available at the inference device, either due to the communication cost of transferring large datasets between such devices, or security or privacy issues. Therefore, this paper seeks to improve the generalization performance of DL-based models for UWA localization via test-time adaptation (TTA)  [^1] Note that in machine learning parlance, test-time adaptation is also called source-free adaptation, where “source” refers to the training domain. However, to avoid confusion with the acoustic source, we use the term test-time adaptation. [17].
To provide a more robust solution than matched field processing (MFP) [1, 18], Chen and Schmidt [7] propose a model-based convolutional neural network (CNN) that outperforms MFP in the presence of sea depth mismatch. It generates the training data using a propagation model with the sound speed profile (SSP) set according to the values measured in a real experiment. This method ignores the parameter shift between the synthetic training data and the real testing data and assumes accurate prior knowledge of the parameters of the test environment, which is rarely available. This motivates further improvement using an adaptation mechanism.
The environmental mismatch problem is known in the machine learning literature as domain shift [19], and has been tackled with approaches including DA or generalization [19, 17], meta learning [20], transfer learning [21, 22], and data augmentation [23]. Transfer learning requires labels for some of the test data, which are usually not available in UWA problems and encourages pursuit of DA methods. Moreover, although DA and transfer learning have been studied for some cases in underwater acoustic localization [21, 24], [11, 25], more extensive investigations into DA are warranted.
Domain adaptation can leverage training data whenever available to learn any underlying domain shift. For UWA localization, inspired by deep subdomain adaptation networks[26], Liu, et al. [11] propose an unsupervised DA method that aims to align the feature distributions of training and test data at every layer of the network. Long, et al. [25] apply an adversarial method based on [13] to extract useful features that are invariant with respect to domain shifts. This approach adds a domain classifier to the original deep model and uses a training loss designed to promote features that confuse this domain classifier in order to achieve domain invariance. These methods, though effective, rely on access to training data. In the absence of training data, TTA is still possible by inferring some useful statistics (features) of the training set from the pre-trained model itself [17].
There is a rich literature on TTA including source hypothesis transfer (SHOT) [17], entropy minimization [27], universal TTA [28], contrastive TTA[29], adaptive adversarial network-based TTA [30], and distribution estimation [31]. To perform TTA, we leverage an implied model uncertainty [32, 33, 34, 35]. Although conformal prediction [36] has recently been adopted for uncertainty quantification in data-driven UWA localization [37], it requires a labeled dataset from the test environment for calibration. To explore the effects of mismatch on model uncertainty without any labeled dataset available at the test time, we adopt model uncertainty using mutual information (MUMI) [34] as a metric. To avoid the computational burden of MUMI during adaptation, in Section III-B, we cast the ranging problem as a classification task that can leverage computationally simpler implied uncertainty quantification methods such as peakwise uncertainty (PU). Our observations in Section III demonstrate that both MUMI and PU increase with an increase in the amount of mismatch, validating their utility.
Environmental mismatches do not affect all samples to the same extent - some samples are affected more and result in a higher uncertainty. For instance, the signal received by the receiver array from a source at a short distance is dominated by the direct path arrivals and first-order reflected raypaths (single reflection from either surface or bottom) [4]. Thus, the depth mismatch affects farther sources more significantly due to the contribution of higher-order bottom reflections received by the array. We study source ranging problems in which there are diverse test samples from different ranges associated with the same source. By partitioning the test samples into certain and uncertain sets based on their PU scores, we can implicitly infer the source power.
Typical data-driven methods in UWA signal processing perform normalization with respect to the received signal strength to obtain source-invariant features [3]. However, this discounts the information in the received signal strength which can aid UWA localization, as demonstrated by some methods [38, 39]. A pre-trained network deprived of the source power information, as discussed in Section III-B, will be confused in the presence of environmental mismatches. To reduce the model confusion, we develop another robust and coarse estimation process based on the received signal strength named the joint source-environment adaptation (JSEA). This leverages the set of certain test samples and enhances the generalization across different environments, while retaining the source-invariance property of the pre-trained model.
The rest of the paper is organized as follows. Section II provides the necessary background on the MFP and CNN approaches for localization. Section III investigates uncertainty quantification in the UWA localization problem. Section IV provides the SHOT approach to adaptation and our proposed method for JSEA. Finally, Section V demonstrates the efficacy of the proposed method on several synthetic and real datasets, followed by discussions on potential extensions and limitations of JSEA in Section VI and some concluding remarks in Section VII.
II Localization Using Sample Covariance Matrices
Notation: Bold letters denote matrices and vectors, $H(\mathbf{y})$ indicates the discrete entropy of a probability mass function (PMF) $\mathbf{y}$, $D_{KL}$ denotes the KL-divergence, $\real{x}$ and $\text{Im}\left(x\right)$ respectively denote the real and imaginary parts of $x$, Superscripts T and H denote respectively the transpose and conjugate transpose of the matrices or vectors. For a set of test points $\mathcal{S}$, the complement is represented by $\mathcal{S}^{\mathsf{c}}=\mathcal{D}_{\text{test}}-\mathcal{S}$ and the cardinality is shown by $|\mathcal{S}|$.
Sample covariance matrices (SCM) are usually used for UWA array-based localization because they are sufficient statistics for the jointly Gaussian signal and noise [40] scenarios. The normalized SCM of an $L$-element array at a frequency $f$ using $P$ snapshots $\mathbf{\tilde{r}}^{(p)}(f)$ is calculated as
$\tilde{\mathbf{C}}(f)=\frac{1}{P}\sum_{p=1}^{P}\mathbf{\tilde{r}}^{(p)}(f)\mathbf{\tilde{r}}^{(p)}(f)^{\mathsf{H}},$ (1)
where $\mathbf{\tilde{r}}^{(p)}(f)=[\tilde{r}^{(p)}_{1}(f),\tilde{r}^{(p)}_{2}(f),...,\tilde{r}^{(p)}_{L}(f)]^{\mathsf{T}}$ and
$\tilde{r}^{(p)}_{l}(f)$ $=\frac{r^{(p)}_{l}(f)}{\sqrt{\sum_{l=1}^{L}|r_{l}^{(p)}(f)|^{2}}},\quad l\in\{1,2,...,L\},$ (2)
where the scalar $r^{(p)}_{l}(f)$ is the complex Fourier coefficient of the $p$-th segment of the received signal on the $l$-th hydrophone at the frequency $f$. For the synthetic data in the simulations, we use the signals generated by the KRAKEN normal-mode model[41] for $\tilde{r}^{(p)}_{l}(f)$, while for the real data, we take the Fourier transform of the recorded temporal waveform and use the Fourier coefficients at the frequency of interest, as explained in Section V-B.
II-A MFP approach
Given that the acoustic source emits a narrowband signal at a frequency $f$, we use the Bartlett processor [42] as an MFP baseline, which is defined as
$\hat{d}=\arg\max_{d}\mathbf{\tilde{r}}(f,d)^{\mathsf{H}}\tilde{\mathbf{C}}(f)\mathbf{\tilde{r}}(f,d),$ (3)
where $\tilde{\mathbf{C}}$ is the normalized SCM of the measured data, and $\mathbf{\tilde{r}}(f,d)$ is the replica field at frequency $f$ generated by a source at range $d$. For each $d$, the corresponding $\mathbf{\tilde{r}}(f,d)$ (of the CNN approach) from the training set is used at inference time.
II-B CNN approach
Localization can be formulated as a regression or a classification problem [7, 43]. Classification models usually assume equidistant labels, as opposed to the regression models that preserve the distance among labels. To address this issue, rather than predicting the true class, we train the classification network to predict a softened label, viz. a PMF, based on the true class. This metric-inspired label softening preserves distances between different classes, makes the model more amenable to the localization problem, and facilitates the proposed uncertainty-based TTA approach. The proposed classification and regression models are both based on a CNN architecture, differing only in the last layer, as in [7].
[…]
$y_{ik}=\frac{\exp(-|k-d^{q}_{i}|/\sigma)}{\sum_{k=1}^{M}\exp(-|k-d^{q}_{i}|/\sigma)},$ (4)
where $M$ is the number of classes and $\sigma$ is a hyperparameter determining the spread of the peak in the soft label. A small $\sigma$ effectively renders the labels one-hot coded while a large $\sigma$ decreases the accuracy by encouraging high-entropy output labels. We empirically conclude that $1\leq\sigma\leq 10$ is appropriate for quantization bins of size $B=100$ m. Fig. 2 shows an example of a soft label with $\sigma=5$.
Figure 1: The CNN classifier includes a feature extraction part followed by a linear classifier that will be frozen during adaptation.
Figure 2: An example of a metric-inspired smoothed label $\mathbf{y}_{i}$ corresponding to $d_{i}=2.9$ km or $d^{q}_{i}=20$. Here, we have used the absolute error as the metric, hence, we have used a truncated exponential as the label. The quantization bins are $100$ m wide.
Therefore, the training dataset is denoted by $\mathcal{D}_{\text{tr}}=\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{N_{\text{tr}}}$ and the test set that does not include labels is denoted by $\mathcal{D}_{\text{test}}=\{\mathbf{x}_{i}\}_{i=1}^{N_{\text{test}}}$. We use the loss function $\mathcal{L}_{tr}=\frac{1}{N_{\text{tr}}}\sum_{i=1}^{N_{\text{tr}}}\mathsf{CE}(\mathbf{y}_{i},\hat{\mathbf{y}}_{i})$ for training, where $\hat{\mathbf{y}}_{i}$ is the network output and $\mathsf{CE}$ is the cross-entropy loss defined by
$\mathsf{CE}(\mathbf{y}_{i},\hat{\mathbf{y}}_{i})\triangleq-\sum_{k=1}^{M}y_{ik}\log\hat{y}_{ik}.$ (5)
The CNN structure used herein is similar to that presented in [7] with a $2$-channel input and $3$ convolutional layers with $6$, $38$, and $40$ channels and kernels of sizes $3$, 55, and $5$, respectively. The output of the feature extraction is a $256$-dimensional vector $\boldsymbol{\phi}_{i}$, which is then fed into the classifier to generate $\hat{\mathbf{y}}_{i}$.
III Model Uncertainty Under Mismatch
For TTA, we first compute the uncertainty of the DL model about its estimations, and then implicitly extract the power of the acoustic source that generated the test data, from those test samples about which the model is most certain. We will discuss two uncertainty quantification methods that can be used with classification or regression models. Although localization naturally fits into a regression paradigm [43], regression models in their standard form only output point estimates. A stochastic regression model (e.g., a dropout network) can produce a PMF output by several realizations of the forward pass [34], which can then be used in MUMI or PU. However, the deterministic classification model already outputs a PMF in its forward pass, which can be used in the PU method.
III-A Model Uncertainty using Mutual Information (MUMI)
There are two different sources of uncertainties in data-driven models [34, 35]: 1) epistemic that corresponds to the lack of training data and 2) aleatoric that corresponds to noisy or inherently ambiguous data. To alleviate the effects of mismatch between the training and test data, we focus on the epistemic uncertainty. Consider a dropout network with the weights vector $\mathbf{w}$. Treating $\mathbf{w}$ as a random variable, the network output for a given test sample $\mathbf{x}_{\mathsf{test}}=\left[\real{\tilde{\mathbf{C}}_{\mathsf{test}}},\text{Im}\left(\tilde{\mathbf{C}}_{\mathsf{test}}\right)\right]$ is also a random variable. If each realization of $\mathbf{w}$ yields a different output for the same input $\mathbf{x}_{\mathsf{test}}$, the model is considered uncertain about the output. MUMI [34] quantifies the amount of information that we would obtain had we received the true label (acoustic source range) $d_{\mathsf{test}}$ for the corresponding $\mathbf{x}_{\mathsf{test}}$. This is represented in terms of the mutual information [44] $I(d_{\mathsf{test}};\mathbf{w}|\mathcal{D}_{\text{tr}},\mathbf{x}_{\mathsf{test}})$, where the conditioning over $\mathcal{D}_{\text{tr}}$ is because we have used the training data to obtain $\mathbf{w}$. Expanding this mutual information yields
$I(d_{\mathsf{test}};\mathbf{w}|\mathcal{D}_{\text{tr}},\mathbf{x}_{\mathsf{test}})$ $=H\left(p(d_{\mathsf{test}}|\mathcal{D}_{\text{tr}},\mathbf{x}_{\mathsf{test}})\right)$
$-\mathbb{E}_{p(\mathbf{w}|\mathcal{D}_{\text{tr}})}\left[H\left(p(d_{\mathsf{test}}|\mathcal{D}_{\text{tr}},\mathbf{w},\mathbf{x}_{\mathsf{test}})\right)\right],$ (6)
where for a given function $h$, $\mathbb{E}_{p(\mathbf{w}|\mathcal{D}_{\text{tr}})}[h(\mathbf{w})]$ denotes the expectation of the random variable $h(\mathbf{w})$ when $\mathbf{w}$ is distributed according to the probability density function $p(\mathbf{w}|\mathcal{D}_{\text{tr}})$. Although computing $\mathbb{E}_{p(\mathbf{w}|\mathcal{D}_{\text{tr}})}[h(\mathbf{w})]$ is intractable in general, a dropout approximating distribution $q_{\theta}(\mathbf{w}|\mathcal{D}_{\text{tr}})$ can efficiently approximate $p(\mathbf{w}|\mathcal{D}_{\text{tr}})$ [34], where $\theta$ represents the parameters of a dropout neural network and each realization of the dropout mask corresponds to a sampling from $q_{\theta}(\mathbf{w}|\mathcal{D}_{\text{tr}})$ that yields a different weight vector $\mathbf{w}_{j}$. Accordingly, the label likelihood $p(d_{\mathsf{test}}|\mathcal{D}_{\text{tr}},\mathbf{x}_{\mathsf{test}})$ in (III-A) can be computed by marginalization over $\mathbf{w}$ and a Monte-Carlo approximation as
$p(d_{\mathsf{test}}|\mathcal{D}_{\text{tr}},\mathbf{x}_{\mathsf{test}})$ $\approx\int p(d_{\mathsf{test}}|\mathbf{w},\mathcal{D}_{\text{tr}},\mathbf{x}_{\mathsf{test}})q_{\theta}(\mathbf{w}|\mathcal{D}_{\text{tr}})d\mathbf{w}$
$\triangleq\frac{1}{J}\sum_{j=1}^{J}\hat{p}(\hat{y}_{j}|\mathbf{w}_{j},\mathcal{D}_{\text{tr}},\mathbf{x}_{\mathsf{test}})$
$=\frac{1}{J}\sum_{j=1}^{J}\hat{p}(\hat{y}_{j}|\mathbf{w}_{j},\mathbf{x}_{\mathsf{test}}),$ (7)
where $\mathbf{w}_{j}\sim q_{\theta}(\mathbf{w}|\mathcal{D}_{\text{tr}}),j=1,...,J$ are weights sampled from the dropout network with $J$ different dropout realizations, $\hat{y}_{j}$ is the network output corresponding to the weight vector $\mathbf{w}_{j}$, and $\hat{p}$ is the empirical likelihood of the model output. Since the training data has been used to obtain $\mathbf{w}$, there is a Markov chain $\mathcal{D}_{\text{tr}}-\mathbf{w}-\hat{y}_{j}$, hence, $\hat{p}(\hat{y}_{j}|\mathcal{D}_{\text{tr}},\mathbf{w}_{j},\mathbf{x}_{\mathsf{test}})=\hat{p}(\hat{y}_{j}|\mathbf{w}_{j},\mathbf{x}_{\mathsf{test}})$. Similarly, $p(d_{\mathsf{test}}|\mathbf{w},\mathcal{D}_{\text{tr}},\mathbf{x}_{\mathsf{test}})$ can be approximated with $\hat{p}(\hat{y}_{j}|\mathbf{w}_{j},\mathbf{x}_{\mathsf{test}})$, which results in
$\mathbb{E}_{p(\mathbf{w}|\mathcal{D}_{\text{tr}})}$ $\left[H\left(p(d_{\mathsf{test}}|\mathcal{D}_{\text{tr}},\mathbf{w},\mathbf{x}_{\mathsf{test}})\right)\right]$
$\approx\frac{1}{J}\sum_{j=1}^{J}H\left(\hat{p}(\hat{y}_{j}|\mathbf{w}_{j},\mathbf{x}_{\mathsf{test}})\right).$ (8)
Substituting (III-A) and (III-A) in (III-A) yields
$I(d_{\mathsf{test}};\mathbf{w}|\mathcal{D}_{\text{tr}},\mathbf{x}_{\mathsf{test}})$ $\approx H\left(\frac{1}{J}\sum_{j=1}^{J}\hat{p}(\hat{y}_{j}|\mathbf{w}_{j},\mathbf{x}_{\mathsf{test}})\right)$
$\quad-\frac{1}{J}\sum_{j=1}^{J}H(\hat{p}(\hat{y}_{j}|\mathbf{w}_{j},\mathbf{x}_{\mathsf{test}})).$
Note that MUMI can be used for both classification and regression networks. While in a classification network the output vector $\hat{\mathbf{y}}_{j}$ is effectively $\hat{p}(\hat{y}_{j}|\mathbf{w}_{j},\mathbf{x}_{\mathsf{test}})$, to obtain the empirical likelihood $\hat{p}(\hat{y}_{j}|\mathbf{w}_{j},\mathbf{x}_{\mathsf{test}})$ in a regression model, we quantize and one-hot encode the output of the regression network, leading to $H(\hat{p}(\hat{y}_{j}|\mathbf{w}_{j},\mathbf{x}_{\mathsf{test}}))=0$. Thus, the MUMI reduces to
$I(d_{\mathsf{test}};\mathbf{w}|\mathcal{D}_{\text{tr}},\mathbf{x}_{\mathsf{test}})\approx H\left(\frac{1}{J}\sum_{j=1}^{J}\hat{p}(\hat{y}_{j}|\mathbf{w}_{j},\mathbf{x}_{\mathsf{test}})\right).$ (9)
Fig. 3 shows that when presented with samples from a mismatched environment, a pre-trained regression model will have a higher MUMI. This shows that the average MUMI increases with the increase in depth mismatch or SSP mismatch. This behavior is similar to the increased MUMI of DL models during adversarial attacks discussed in [34]. Also, MUMI is seen to be a monotonically increasing function of the amount of mismatch, validating its use as a metric of uncertainty.
(a) Uncertainty under SSP mismatch.
(b) Uncertainty under depth mismatch.
(c) Average uncertainty over all ranges considered under SSP mismatch.
(d) Average uncertainty over all ranges considered under depth mismatch.
Figure 3: MUMI uncertainty of a regression dropout network reveals the mismatch between the training and test environments. These are using the SWellEx-96-inspired synthetic data, consisting of a shallow source with a monotone signal at $f=109$ Hz and all of the results are averaged over $100$ independent realizations of an additive white Gaussian noise (AWGN) channel. Top row shows that the uncertainty is different for each source range and the bottom row shows that, on average, both the uncertainty and the root-mean-squared-error (RMSE) increase with the amount of mismatch.
III-B Peakwise Uncertainty (PU)
While MUMI is theoretically motivated, the MC-dropout requires $J$ realizations of the dropout layers to obtain $\{\hat{p}(\hat{y}_{j}|\mathbf{w}_{j},\mathbf{x}_{\mathsf{test}})\}_{j=1}^{J}$ which is not feasible for underwater low-power applications. Here, we propose a simpler method to assess uncertainty in the classification approach tailored to the localization problem. This method counts the number of significant peaks in the output PMF, and is based on the intuition that the model is certain about the corresponding result if there is only one significant peak, whereas in cases where noise or environmental mismatch dominate, there would not be one significant peak only.Following this approach, we define a “significant peak” as a peak that is either the only one or is larger than the next-largest peak by a factor of at least $Q$. The hyperparameter $Q$ should be selected to draw a trade-off between the desired size of $\mathcal{S}$ and the minimum confidence desired in pseudo labels selected. Thus
[…]
where $\hat{\mathbf{y}}$ is the classification network output probability vector. Throughout the paper, we use $Q=10$. The average PU indicates the percentage of samples about which the model is uncertain and is defined as
$APU[\%]=100\frac{\sum_{i=1}^{N}PU(\mathbf{x}_{i})}{N}.$
Fig. 4 shows that the APU increases as the amount of mismatch increases.
(a) Uncertainty under SSP mismatch.
(b) Uncertainty under depth mismatch.
Figure 4: Variation of the average peakwise uncertainty with the amount of environmental mismatch in (a) depth and (b) SSP.
Peak counting is not commonly used for uncertainty quantification. Nevertheless, APU shows a similar variation as MUMI in the presence of mismatches. This motivates interpreting PU as an implied uncertainty. As illustrated in the example in Fig. 5, the largest peak in the output of the pre-trained model is not necessarily aligned with the true label. The JSEA algorithm introduced in the next section, determines which peak to select.
Figure 5: An example of a predicted label in a test environment with a different SSP than the training environment. The quantization bins are $100$ m wide. Observe that the pre-trained model prediction (the label with maximum probability) is far from the ground-truth. However, there is another peak in the output closer to the ground-truth. After JSEA-based adaptation, the output shows one clear peak at the true label, showing how JSEA corrects and adapts the network to give correct labels in the presence of environmental mismatch.
IV Test Time Adaptation
Our proposed TTA method, similar to SHOT [17], assumes that the DL-based model consists of a domain-dependent feature extractor and a domain-independent classifier (hypothesis). During inference, while freezing the classifier, SHOT fine-tunes the feature extractor to produce confident outputs and prevents it from a degenerate solution using an information maximization loss [17] that promotes diverse output labels. Although this procedure works well for one-hot coded labels, it must be adjusted for ranging tasks, where labels are not equidistant.
IV-A Environmental Adaptation Using SHOT
When presented with data from an environment that is markedly different from the one on which it was trained, the network makes uncertain predictions, i.e., it is likely to be more confused between different classes, which leads to inaccurate predictions. During inference, the absence of test labels makes fully-supervised adaptation impossible. Nevertheless, encouraging the model to more confidently preserve its predictions on samples that match to the training data, helps it overcome slight mismatches. To this end, we select a self-supervising subset $\mathcal{S}$ of the samples and use the pre-trained network’s output on these samples as their pseudo-labels, for which we use a loss term $\mathsf{CE}(\mathbf{y}_{i}^{\mathsf{pseudo}},\hat{\mathbf{y}}_{i})$. Here, $\mathbf{y}_{i}^{\mathsf{pseudo}}$ is the softened label based on the output of the pre-trained classifier. However, this loss and approach can create a degenerate situation where the network only predicts classes that exist in $\mathcal{S}$ [17]. Adding a regularizing loss term to maximize the entropy of the outputs average $\bar{\mathbf{y}}=\frac{1}{N_{\text{test}}}\sum_{i=1}^{N_{\text{test}}}\hat{\mathbf{y}}_{i}$ can prevent this situation.
At test time, we freeze the pre-trained classifier as depicted in Fig. 1, which predicts the label $\hat{\mathbf{y}}_{i}=[\hat{y}_{i1},\hat{y}_{i2},...,\hat{y}_{iM}]$ by applying the Softmax() operation to $\boldsymbol{\omega}_{k}^{\mathsf{T}}\boldsymbol{\phi}_{i}$, defined as,
$\hat{y}_{ik}=\frac{\exp(\boldsymbol{\omega}_{k}^{\mathsf{T}}\boldsymbol{\phi}_{i})}{\sum_{k=1}^{M}\exp(\boldsymbol{\omega}_{k}^{\mathsf{T}}\boldsymbol{\phi}_{i})},$ (11)
where $\boldsymbol{\omega}_{k}$ is the classification weight vector corresponding to the $k$-th class. Note that if the feature vector $\boldsymbol{\phi}_{i}$ is highly aligned with $\boldsymbol{\omega}_{k_{0}}$ for only one $k_{0}$, then the network predicts a unimodal soft label $\hat{\mathbf{y}}_{i}$, as opposed to situations where the network generates a multimodal output and is not certain about the predicted label. Inspired by this observation, we construct the self-supervising subset $\mathcal{S}$ according to
$\mathcal{S}=\{i\;:1\leq i\leq N_{\text{test}}\,\,\text{s.t.}\,\,PU(\mathbf{x}_{i})=0\}.$ (12)
To perform the adaptation, we minimize the SHOT loss defined by
$\mathcal{L}_{\text{SHOT}}$ $=-H\left(\bar{\mathbf{y}}\right)+\frac{\beta}{|\mathcal{S}|}\sum_{i\in\mathcal{S}}\mathsf{CE}(\mathbf{y}_{i}^{\mathsf{pseudo}},\hat{\mathbf{y}}_{i}),$ (13)
where $\beta$ is a positive hyperparameter. We then employ the Adam optimizer [45] to minimize $\mathcal{L}_{\text{SHOT}}$ with a step size of $\mu_{\text{SHOT}}$ to update the CNN feature extractor. The cross-entropy term encourages the network to remain certain about its existing confident predictions, while increasing $H\left(\bar{\mathbf{y}}\right)$ prevents it from assigning all inputs to the same output by encouraging diverse outputs [17]. Fig. 6 provides a flowchart for adapting the model using the SHOT algorithm.
Figure 6: A flowchart for the SHOT approach for adapting the localization model. There are three major steps in SHOT: First, the CNN model estimates the labels; then the set of certain samples $\mathcal{S}$ is determined. For the samples in $\mathcal{S}$, the CNN estimates are considered the final range estimates (pseudo-labels). Then, the CNN is retrained using the pseudo-labeled samples in $\mathcal{S}$. Finally, for each of the uncertain samples in $\mathcal{S}^{c}$, the final range estimate is obtained from the retrained CNN.
Although effective at rectifying errors due to small mismatches, SHOT does not directly enhance the estimates for samples in $\mathcal{S}^{c}$ and does not focus on removing the model’s confusions. To alleviate this, we now introduce another adaptation mechanism that explicitly selects pseudo-labels for the samples in $\mathcal{S}^{c}$.
IV-B Joint Source-Environment Adaptation (JSEA)
While TTA methods in the machine learning literature focus on domain adaptation, they do not consider or distinguish the source that generates data and the environment where data is generated. In UWA, however, there is a clear distinction between these two factors. To improve the generalization performance of data-driven methods in UWA signal processing, a necessary step is normalization with respect to the source, which effectively yields source powerinvariant features and better generalization across different sources. Accordingly, array-based localization usually involves normalization by the total received signal power at the array elements. However, by doing this, one loses valuable information on the source power that could be leveraged to improve localization. For example, as shown in Fig. 7, the array SNR at the given frequency can provide a range estimate that (i) is coarse, due to the local fluctuations in the signal power level, but (ii) is robust because the general trend is that higher SNR is associated with a closer source, regardless of the environment. To exploit the distinction between the source and the environment, we implicitly estimate the test source power from a limited set of reliable samples (samples with a unimodal output PMF). We then develop a signal strength-based approach to estimate the range of more difficult samples (samples with a multimodal output PMF). This results in combining the DL-based and signal strength-based approaches in a Bayesian manner to fully use the information in an efficient manner.
Figure 7: The received SNR (blue) on the vertical linear array at frequency $f=130$ Hz, emitted from the deep source in the event S5 of the SWellEx-96 experiment [46], and source range from the receiver (red). The gaps in the SNR plot indicate some time intervals during which transmission was stopped and the array recorded only the ambient noise. These were used for the SNR calculation.
The method is detailed as follows. Given that $\mathbf{C}=\psi\;\tilde{\mathbf{C}}$, where $\psi=\sum_{l=1}^{L}|r_{l}^{(p)}(f)|^{2}$ (which is an accurate definition of received signal power for $P=1$ snapshot case and an approximate estimate for $P>1$), and assuming that $\psi$ and $\tilde{\mathbf{C}}$ are independent, by Bayes’ theorem
$p(d|\mathbf{C})$ $=p(d|\tilde{\mathbf{C}},\psi)=\frac{p(\tilde{\mathbf{C}},\psi|d)p(d)}{p(\tilde{\mathbf{C}})p(\psi)}$
$=\frac{p(\tilde{\mathbf{C}}|\psi,d)p(\psi|d)p(d)}{p(\tilde{\mathbf{C}})p(\psi)}=\frac{p(d|\tilde{\mathbf{C}})p(d|\psi)}{p(d)}.$ (14)
[…]
$\hat{d}_{i}=\arg\max_{d\in\mathcal{P}(\hat{\mathbf{y}}_{i})}p(d|\psi=\psi_{i}).$ (15)
In a homogeneous boundary-less acoustic medium, the received power (on a single hydrophone) from a point source at distance $d$ is proportional to $1/d$. In ocean waveguides, however, the received power fluctuates around a decreasing function $\psi_{0}(d)$; the fluctuations are affected by the ocean variability and ambient noise as well, leading to the low resolution of the estimates based on $p(d|\psi)$. Assuming that $\psi(d)$ is normally distributed with the mean $\psi_{0}(d)$ and considering a non-informative prior distribution on $\psi$, (15) reduces to
$\hat{d}_{i}=\arg\min_{d\in\mathcal{P}(\hat{\mathbf{y}}_{i})}(\psi_{i}-\psi_{0}(d))^{2}.$ (16)
Since $\psi_{0}$ is unknown, we approximate $\psi_{0}(d)$ for each $d\in\mathcal{P}(\hat{\mathbf{y}}_{i})$ using the samples in $\mathcal{S}$ with [12] $\hat{\psi}_{0}(d)=\frac{1}{|\mathcal{S}_{\delta}(d)|}\sum_{j\in\mathcal{S}_{\delta}(d)}\psi_{j}$, where $\mathcal{S}_{\delta}(d)=\{j\in\mathcal{S};|d-\hat{d}_{j}|\leq\delta\}$, and $\psi_{j}$ is the received power of the $j$-th sample. Therefore,
$\hat{d}_{i}=\arg\min_{d\in\mathcal{P}(\hat{\mathbf{y}}_{i})}\left(\psi_{i}-\frac{1}{|\mathcal{S}_{\delta}(d)|}\sum_{j\in\mathcal{S}_{\delta}(d)}\psi_{j}\right)^{2}.$ (17)
By defining the pseudo-labels for $\mathbf{x}_{i}$ as $\mathbf{y}_{i}^{\mathsf{pseudo}}\triangleq\eta(\hat{d}_{i})$, we have achieved the range estimates for all given test samples. If the model is to be used for another source in the same environment, one should also adapt the feature extractor by minimizing the JSEA loss defined by $\mathcal{L}_{\text{JSEA}}=\sum_{i=1}^{N_{\text{test}}}\mathsf{CE}(\mathbf{y}_{i}^{\mathsf{pseudo}},\hat{\mathbf{y}}_{i})$. Fig. 8 provides a flowchart for the JSEA algorithm.
While SHOT defines the pseudo-labels only for the samples in $\mathcal{S}$, JSEA defines pseudo-labels for all test samples. In both SHOT and JSEA, pseudo-labels for certain samples are the final range estimates. However, to obtain the final range estimates for uncertain samples, SHOT relies on re-training the network for a few iterations using the pseudo-labeled certain samples, while JSEA relies on the peak selection process based on the received signal power.
Figure 8: A flowchart for JSEA. There are two major steps in JSEA: First, the CNN model estimates the labels; then the set of certain samples $\mathcal{S}$ is determined. For the samples in $\mathcal{S}$, the CNN estimates are considered the final range estimates. Finally, for each of the uncertain samples in $\mathcal{S}^{c}$, the final range estimate is selected from the set of peak locations, according to (17).
V Evaluation
V-A Synthetic data
We first consider data generated using KRAKEN [41], where the environmental parameters are set according to the SWellEx-96 [42, 10] experiment. As depicted in Fig. 9, the training data consist of source ranges between $0.85$ km and $9.05$ km with $10$ m increments (thus, $821$ distinct samples), and a $21$-element vertical linear array placed at depths from $94.125$ m to $212.25$ m to mimic the SWellEx-96 data. The source depth is assumed to be $9$ m and transmits a monotone signal at a frequency of $f=109$ Hz. We consider a flat bathymetry at depth $216.5$ m in our simulations and used the average SSP from the SWellEx-96 measurements. The bottom constitutes of $3$ layers whose depths, densities, and attenuation match the SWellEx-96 environment.
Figure 9: KRAKEN environment based on the SWellEx-96 experiment. The receiver array is a 21-element array depicted in red. The uniformly sampled source ranges that are used for training are shown with black dots. Also the SSP, which is the average SSP from the SWellEx-96 experiment, is shown with the blue graph.
To generate a test environment that has a sound-speed $c(z)$ at depth $z$, which is mismatched with that of the training environment, $c_{0}(z)$, we perturb $c_{0}(z)$ using a constant gradient difference of $\Delta c$ (m/s)/m, represented as $c(z)=c_{0}(z)+\frac{\Delta c}{216.5}(z-216.5)$, where $0\leq z\leq 216.5$ m. We generate data for $500$ random source ranges between $D_{\min}=900$ m and $D_{\max}=9$ km, and divide the region into $100$ m-long intervals, each represented by its mid point. Consequently, the labels and classifier outputs are 82-dimensional (i.e., $M=82$). In addition, to emulate the real ocean ambient noise during test time, we add a randomly selected noise segment $w$ from the KAM11 experiment [47] (scaled to have a desired signal-to-noise ratio in the data) to the generated signal for each hydrophone. Since the proposed method takes a batch of test data, we use the average array SNR of the batch, defined as
$\text{SNR [dB]}=10\log_{10}\frac{\sum_{i=1}^{N_{\text{test}}}\sum_{l=1}^{L}|r_{i,l}|^{2}}{N_{\text{test}}L\psi_{w}},$
where $|r_{i,l}|^{2}$ is the received signal power at the $l$-th hydrophone of the $i$-th sample and $\psi_{w}$ is the noise power spectral density at the frequency $f=109$ Hz.
We evaluate models using MAE and probability of credible localization (PCL) [21] defined as
$\text{PCL}(\zeta)$ $=\frac{100}{N_{\text{test}}}\sum_{i=1}^{N_{\text{test}}}\mathbf{1}_{|d_{i}-\hat{d}_{i}|\leq\zeta d_{i}},$
MAE $=\frac{1}{N_{\text{test}}}\sum_{i=1}^{N_{\text{test}}}|d_{i}-\hat{d}_{i}|,$
where $d_{i}$ and $\hat{d}_{i}$ are respectively the true and estimated distances of the $i$-th sample. Probability of credible localization can be interpreted as the probability that the range estimate lies within a certain percentage bound of the true range, defined here as $\zeta=10\%$.
The methods evaluated include:
  • O-MFP (Oracle-MFP): The MFP approach tuned to the test environment. This method is presented just for reference and is expected to outperform all other methods because it has access to the test environment parameters.
  • M-MFP (mismatched MFP): The MFP approach tuned to the training environment. This method is assumed to have access to the training data.
  • CNN-c: A CNN-based model for classification, trained once on the training set, and not adapted to account for mismatch.
  • SHOT: The CNN-c equipped with the adaptation mechanism in (13), with $\mu_{\text{SHOT}}=5\times 10^{-6}$ and $\beta=1$.
  • JSEA-c: The CNN-c equipped with the pseudo-label estimator in (17), where $\delta=500$ m.
  • CNN-r: A CNN-based model for regression.
  • JSEA-r: The CNN-r followed by an MC-dropout mechanism to generate output PMFs and equipped with the pseudo-label estimator in (17), where $\delta=500$ m.
To select the label smoothing hyperparameter $\sigma$, we compared the performance of the trained model on the validation data for different values of $\sigma$ and found that $\sigma=2$ works well. We also use a learning rate of $10^{-4}$ and a batch size of $128$ for the Adam optimizer. We randomly split the set of uniformly sampled ranges into $82\%$ training and $18\%$ validation sets. To prevent over-fitting and converge better to the minimum, we adopt a decreasing learning rate - we multiply the learning rate by $0.1$ whenever there is no reduction in the validation error after $75$ iterations, and we stop training if there is no reduction in the validation error after $125$ iterations [7]. To enhance robustness against noise, after training the network with clean training data, we continue training with noisy data, where the inputs are contaminated by white Gaussian noise with an SNR in $\{2,4,6,8,10,12,14,16\}$ dB. The test results are averaged over $100$ noise realizations.
Fig. 10 shows the superiority of the JSEA-c and CNN-based classification methods over M-MFP when the training and testing environments have a depth mismatch of $\Delta D=4$ m (i.e, the ocean depth in the test environment is $220.5$ m), and there is a mismatch in the noise statistics of the test data compared to the training data as well. The proposed JSEA-c outperforms other methods including SHOT, however, its improvement highly depends on the success of the pre-trained model in generating enough certain labels. Although, JSEA-r can successfully improve the regression results as seen in Fig. 10, we will not include the regression models in the following experiments due to their inferior performance compared to the classifier-based approaches. In this scenario, the M-MFP shows inferior performance with respect to classification methods, in the presence of depth mismatch.
(a) PCL(10%)
(b) MAE
Figure 10: Performance versus SNR using noise data from KAM11 experiment and with $\Delta D=4$ m for the shallow source with the monotone signal at $f=109$ Hz, evaluated with metrics: (a) PCL and (b) MAE. While the O-MFP provides a performance bound as it has access to training data generated in an environment matched to the test environment, the M-MFP shows the performance degradation of the MFP method when provided with the mismatched data.
Fig. 11 shows the performance comparison of the methods with variation in the degree of SSP mismatch. In the current scenario, mismatch in SSP gradient is assessed, by varying the gradient difference $\Delta c$ between training and test data. As seen in Fig. 11, the M-MFP outperforms the DL models; however, it benefits from using training data during inference, which is not available to the other methods. While SHOT can improve the CNN-c results, JSEA-c has the best performance among the DL methods.
(a) PCL(10%) under mismatch in SSP.
(b) MAE under mismatch in SSP.
Figure 11: Performance against variation in SSP mismatch at an SNR of $15$ dB.
Next we consider the depth mismatch and evaluate the performance of the algorithms for $\Delta D\in\{0,2,4,6\}$ m/s. As seen in Fig. 12, when the test and training environments match (i.e., $\Delta D=0$), the MFP (in this case, the M-MFP is not mismatched, hence it is the same as O-MFP) performs the best. However, depth mismatches lead to the M-MFP’s performance deteriorating, whereas the performance of the CNN-based methods is more robust to this mismatch. Thus, using CNN methods can impart some robustness to localization performance under depth mismatch.
(a) PCL(10%) under mismatch in depth.
(b) MAE under mismatch in depth.
Figure 12: Performance against variation in magnitude of depth mismatch at an SNR of $15$ dB.
To further investigate the effect of environmental mismatches, we evaluate the performance against mismatches in bathymetric shape/fluctuations, as shown in Fig. 13. While the training environment includes a flat bathymetry, the test environment contains a seamount. Fig. 14 shows the MAE and PCL results versus the seamount height. The results in Fig. 14 verify the superior performance of the JSEA-c in terms of both MAE and PCL for this scenario.
(a) Bathymetry of the training environment, inspired by the SWellEx-96 environment.
(b) Mismatched bathymetry used for testing. There is a seamount structure at 1 km distance from the source, with a height of $h$ meters.
Figure 13: Bathymetry mismatch used to evaluate the performance of the JSEA method.
(a) PCL(10%) under mismatch in bathymetry.
(b) MAE under mismatch in bathymetry.
Figure 14: Performance against variation in bathymetry at an SNR of $15$ dB.
Mismatch in bottom sediment type is another type of environmental mismatch that commonly arises in localization problems. Each sediment type varies in density, sound speed, and attenuation coefficient, which affect the acoustic wave propagation through bottom interactions. For the sake of demonstrating robustness to mismatch, we consider only fluid sediment types here, and do not consider shear speed or attenuation. To evaluate the performance against mismatches in sediment type, we generate training data according to the sediment type provided in Fig.13(a) and evaluate the performance on test data generated in environments with sediment types defined according to Table I [48]. Fig. 15 shows the MAE and PCL results versus the sediment type. These results verify the superior performance of the JSEA-c in terms of both MAE and PCL for this scenario.
TABLE I: Different sediment types [48]
training sediment clay silt sand gravel moraine
density $\rho\;\;[g/cm^{3}]$ 1.76 1.5 1.7 1.9 2 2.1
minimum sound speed $c_{\text{min}}[m/s]$ 1572.37 1500 1575 1650 1800 1950
maximum sound speed $c_{\text{max}}[m/s]$ 1593.02 1520 1595 1670 1820 1970
attenuation $\alpha$ [dB/ (km Hz)] 0.2 0.2 1 0.8 0.6 0.4
(a) PCL(10%) under mismatch in sediment type.
(b) MAE under mismatch in sediment type.
Figure 15: Performance against mismatch with various sediment types at an SNR of $15$ dB.
Overall, we observe that using TTA can enhance the performance of the CNN-based approach by allowing better adaptation to the test environment and introducing more robustness to mismatch. This is demonstrated by the improvement in SHOT and JSEA in comparison to CNN. Using the received signal power information further improves the performance, as seen in the improvement in JSEA as compared to SHOT.
V-B SWellEx-96 Data
We use data from the event S5 of the public dataset SWellEx-96 experiment [46, 2, 49]. In this event, two sources, a shallow source at a depth of $9$ m and a deep source at a depth of $54$ m, were towed by a ship moving in a straight line as shown in Fig. 16. The whole event took $75$ minutes, during which each source transmitted a set of frequencies with a specific power level, recorded by several arrays at the sampling frequency of $F_{s}=1500$ Hz. We use the recordings from the $21$-element ($L=21$) vertical line array (VLA) and provide the results for the shallow source.
To obtain the Fourier transform and the corresponding SCM $\mathbf{C}$ used for localization, we use a $3$-second-long signal divided into $P=5$ segments, each of length $1$ second and a $50\%$ overlap, tapered by a Kaiser window with $\beta=9.24$. We obtain the true source range by linear interpolation between the GPS recordings provided assuming the ship followed a straight-line path between GPS updates.
Figure 16: The ship track for the SWellEx-96 S5 event [46]. The localization problem considered involves estimating the distance of the ship to the VLA receiver.
Fig. 17 shows the ranging results of M-MFP, CNN-c, and JSEA-c on the real-world narrowband data at frequency $f=109$ Hz. The results in Fig. 17(a) suggest that M-MFP performs poorly in this scenario, as it leans towards overestimating the ranges for the samples around $t=60$ min which are closest to the VLA and have a relatively higher SNR. This problem is absent in the CNN-c approach as observed in Fig. 17(b). Nevertheless, both M-MFP and CNN-c tend to deviate significantly from the true ranges in the low SNR regions, i.e., larger source ranges ($t\leq 40$ min). While this problem is still present in JSEA-c, it is considerably less severe, yielding an overall better MAE and PCL, as shown in Table II.
To understand the effect of the JSEA-c approach on reducing the error, we show the partitioning of the samples into certain and uncertain sets based on the PU scores, in Fig. 4. This figure denotes that most of the uncertain samples are in the low SNR region, where the M-MFP and CNN-c fail. This means that the JSEA-c approach relies on the certain samples (mainly the correctly estimated samples which tend to be from the high SNR region) to rectify the incorrect outputs of the CNN-c. This might also be, in part, due to the fact the samples around $t=60$ min are not only higher in SNR, but also their environment is matched the most to the training environment, as the depth we use for training environment is $216.5$ m, which is the depth at the VLA location.
(a) M-MFP.
(b) CNN-c.
(c) JSEA-c.
Figure 17: Localization results for SWellEx-96 data. The blue curve indicates the ground-truth values of range at each time. The orange dots in each plot, indicate the corresponding estimated ranges. The gray cone around the blue curve indicates an interval from $0.9$ to $1.1$ of the true range. Observe how with JSEA, the predicted ranges are concentrated closer to the true ranges.
TABLE II: Localization performance for SWellEx-96 data.
M-MFP CNN-c JSEA-c
MAE [km] 2.387 1.746 1.430
PCL ($20\%$) 31.78 38.89 42.67
(a) Certain examples.
(b) Uncertain examples.
Figure 18: Partitioning of the CNN-c predictions for SWellEx-96 data into certain and uncertain examples using PU metric. The solid blue line indicates the ground-truth. The left plot shows certain examples and the right plot shows uncertain examples.
VI Discussion, Extensions, And Limitations of JSEA
Whereas the empirical evaluations in the previous section verify the efficacy of JSEA in several different scenarios, it is a sophisticated method with some inherent limitations. This section addresses hyperparameter selection, JSEA limitations and possible algorithmic extensions, and complexity comparison with other methods.
VI-A Hyperparameter Selection
Successful operation of JSEA relies on reasonable selection of the hyper-parameters. The hyper-parameters $\sigma$ and $B$ are related to converting the regression problem to a classification one, and can be selected by evaluating the performance on a validation set, and adjusting their values according to the following guidelines. Moreover, the adaptation hyper-parameters $\delta$ and $Q$ can be determined based on the desired performance and any prior knowledge on the degree of the mismatch between training and test environments, according to the following discussion. In the following, we explain the effect of each of these four hyper-parameters.
  • $\sigma$: A large $\sigma$ leads to ground-truth labels that are heavy-tailed and closer to a uniform distribution and cannot indicate the distinction between different classes clearly. This in turn slows the training due to weak supervisory signal and can lead to an inferior accuracy. On the other hand, a small $\sigma$ leads to ground-truth labels that are light-tailed and closer to a point mass (spike), rendering labels one-hot-coded, which can exacerbate the overfitting of the model by removing the distance information between different classes. Hence the $\sigma$ value is chosen to balance the tradeoff between these two factors.
  • $B$: A larger $B$ leads to lower precision (a coarser quantization) in localization which bounds the best localization performance possible, but also yields a less complex model with fewer output classes which makes classification easier and enhances performance. So in practice, $B$ should be determined according to this tradeoff between the precision required and the desired model complexity. Based on this tradeoff, and also following Chen and Schmidt [1], who used a similar CNN-based architecture for source ranging using the SWellEx-96 dataset, we select $B=100$ m.
  • $Q$: The hyperparameter $Q$ adjusts the sensitivity of the algorithm to mismatches; the larger $Q$ is, the more sensitive JSEA becomes to mismatches. As such, $Q$ should be selected to draw a trade-off between the desired size of $\mathcal{S}$ and the minimum confidence desired in pseudo-labels selected. Determining this trade-off is naturally tied to $\sigma$ and $\delta$. We empirically observed that $Q=10$ leads to a reasonable trade-off.
  • $\delta$: This hyper-parameter should be small enough to be able to distinguish among the peaks to rectify the labels of uncertain samples, i.e.,
    $\forall j\in\mathcal{S}^{c}:\;\;\forall d_{1},d_{2}\in\mathcal{P}(\hat{\mathbf{y}}_{j}),\;\;\text{if}\;\;d_{1}\neq d_{2},\;\;\mathcal{S}_{\delta}(d_{1})\neq\mathcal{S}_{\delta}(d_{2}),$
    and it should be large enough that
    $\forall j\in\mathcal{S}^{c}:\;\;\forall d\in\mathcal{P}(\hat{\mathbf{y}}_{j}),\;\;\mathcal{S}_{\delta}(d)\neq\phi.$
    If such a $\delta$ cannot be found for a sample $j$, one can either accept the largest output peak as the final estimate, or declare the sample as an uncertain one that cannot be rectified using JSEA. We have used $\delta=500$ m for all test samples as this value satisfied these conditions.
VI-B Complexity Comparison
Underwater applications typically require power- and electronic storage-efficient algorithmic designs. Here, we provide a complexity comparison between different methods that were discussed in the paper. Although the calculations are not based on hardware-optimized algorithms, they provide a reasonable baseline for comparison. The complexity analysis is provided in Appendix A. Table III summarizes the complexity and run-time comparison between different methods, when implemented on an NVIDIA RTX 6000 Ada Generation GPU device using CUDA 12.2.
TABLE III: Memory and Computational Complexity Comparison
Model Computational Complexity Memory Complexity Run Time $[ms]$
MFP $O(N_{\text{tr}}L^{2})$ $O(N_{\text{tr}}L)$ $0.4666$
CNN-c $O(N_{\phi}L^{2}+MN_{\phi})$ $O(N_{\phi}L^{2}+MN_{\phi})$ $0.0014$
SHOT $O(N_{\phi}L^{2}+MN_{\phi}+N_{\text{itr}}N_{\phi}L^{2})$ $O(N_{\text{test}}L+N_{\phi}L^{2}+MN_{\phi})$ $1.5225$
JSEA-c $O(N_{\phi}L^{2}+MN_{\phi}+N_{\text{test}})$ $O(N_{\text{test}}L+N_{\phi}L^{2}+MN_{\phi})$ $0.2353$
VI-C Extensions And Limitations
The proposed JSEA mechanism relies on access to a sufficient number of test samples as well as the ability of the CNN method to classify a sufficient number of them with certainty. Determining the minimum number of certain samples required for successful operation of JSEA is left for future research. Furthermore, variations of JSEA can be developed that use other assumptions for $p(d|\psi)$ in (15) and correspondingly may lead to more efficient methods for source power estimation from the samples in $\mathcal{S}$.
In the presence of more than one acoustic source at the intended frequency, MFP and CNN-based methods may fail in localization. In such scenarios, the JSEA method might show spurious peaks even in an environment that closely matches the training environment, i.e., JSEA reveals a mismatch in the situation (multi-source situation during inference versus the single-source situation during the training). Nevertheless, JSEA is not able to rectify such errors, as it requires samples from the same source at different ranges to implicitly estimate the source power.
The simulations in this paper verify the efficacy of JSEA for narrowband sources. For broadband source localization, there are several approaches to exploit the information contained in different frequency bands. One approach is using the SCMs obtained for multiple frequencies as inputs of size $2N_{F}\times L\times L$ to the CNN model, where $N_{F}$ is the number of frequencies. In this scenario, JSEA can be used in the same manner it is used the narrowband case. However, another approach is to train a different sub-model for each frequency and then aggregate their estimates to obtain the source range. In this scenario, we can either apply JSEA to each sub-model first and then aggregate the results, or first aggregate the results (e.g., by averaging the output PMFs from all sub-models) and then use JSEA as in the narrowband scenario. The efficacy of this approach remains to be explored in future work.
VII Conclusion
In this paper, we tackled test-time adaptation of a deep-learning based localization algorithm to environmental mismatch between training and test data. We considered different forms of mismatch including that in water depth, bathymetric shape and sediment type. We showed that an appropriate model uncertainty such as MUMI has the potential to reveal the mismatch between the training and test environments. Since regression models equipped with MUMI are computationally complex, we resort to a classification paradigm, using a special label softening to tailor the method to localization. By using classification for localization, the model outputs reveal the peakwise uncertainty, which we then use to enhance the localization based on the source power.
We showed that while SHOT can slightly improve the performance by encouraging the network to make its predictions less uncertain on the samples that are similar to the training set, JSEA can further improve the UWA localization performance on the samples that suffer more from the environmental mismatch. To be successful, the proposed JSEA method requires a pre-trained model that can correctly classify part of the test data and do so with certainty.
This paper shows that by suitable domain adaptation, one can enhance localization robustness in data-driven methods. With more prior knowledge about domain shifts between the training and testing environments and source location distribution, one may be able to improve the performance by using more effective methods for selecting $\mathcal{S}$ even without access to training data. However, the performance improvement will be limited in the absence of such knowledge. Future potential directions on this front include extending the method to broadband sources or multiple sources, investigating other TTA methods currently used for other machine learning problems and developing other methods for uncertainty quantification in the stated UWA problem.
Appendix A Complexity Analysis
Figure 19: A typical convolutional layer with $H_{\text{in}}$ input channels and $H_{\text{out}}$ output channels.
Fig. 19 shows the size of tensors involved in a typical convolutional layer, where there are $H_{\text{in}}$ input channels, there are $H_{\text{out}}$ output channels, and the input and output dimensions are both $L\times L$ due to to padding. According to this figure, the number of multiplications in a convolutional layer is $H_{\text{in}}H_{\text{out}}L^{2}k^{2}$ and the number of additions is $H_{\text{in}}H_{\text{out}}L^{2}k^{2}$, resulting in a computational complexity of $2H_{\text{in}}H_{\text{out}}L^{2}k^{2}$. This yields a computational complexity of $O(L^{2})$ and a memory complexity of $H_{\text{in}}H_{\text{out}}L^{2}$. Moreover, according to Fig. 19, each convolutional layer require $H_{\text{in}}H_{\text{out}}k^{2}$ floating point numbers to be saved, while each linear layer requires a matrix of size input-size $\times$ output-size to be saved. Also, each layer requires an activation memory to hold the input and output tensors. Table IV summarizes the computational and memory complexities of each layer during a forward pass. In addition, updating the network weights using samples from the test environment involves backpropagation. Each backpropagation iteration requires the same amount of memory that a forward pass needs, to save the gradients with respect to inputs and weights. However, the computational complexity of a backpropagation pass is twice that of a forward pass, as backpropagation involves computing gradients with respect to not only the weights, but also inputs.
TABLE IV: CNN Architecture Complexity Analysis
Layer Type $H_{in}$ $H_{out}$ Input Size Output Size
Computational
Complexity
Memory
Complexity
1 conv (k=3) 2 6 $\times L L×L$$\times L$ $216L^{2}$ $2L^{2}+6L^{2}+108$
2 conv (k=5) 6 38 $\times L L×L$$\times L$ $11{,}400L^{2}$ $38L^{2}+5{,}700$
3 conv (k=5) 38 40 $\times L L×L$$\times L$ $76{,}000L^{2}$ $40L^{2}+38{,}000$
4 linear 11 11 $40L^{2}$ $N_{\phi}$ $80N_{\phi}L^{2}$ $40L^{2}N_{\phi}$
5 linear 11 11 $N_{\phi}$ $M$ $2N_{\phi}M$ $MN_{\phi}$
The JSEA methods refines the outputs of the CNN to achieve the final estimates for the source ranges. While the aforementioned complexity analysis applies to any general CNN architecture, the JSEA refinement step requires saving the estimates and energy levels of the received signals for a subset of the test samples (certain samples). Therefore, this refinement step increases the memory complexity by $O(N_{\text{test}})$ floating point numbers. Furthermore, peak finding algorithm may be performed with $O(MW)$ complexity, where $W$ is the size of the sliding window over which we find the local maximum and $M$ is the number of classes. Also, JSEA involves computing the received signal power, which adds to the complexity by $O(M)$ per sample. Finally, for an uncertain sample with $N_{P}$ significant peaks, we need to construct $\mathcal{S}_{\delta}(d)\subseteq\mathcal{S}$ for each peak and compare the average received power over that set to that of the uncertain sample, which can be done with a complexity of $O(N_{P}|\mathcal{S}|)$. Note that while JSEA refinement provides final range estimates, fine-tuning the model (model adaptation) using the pseudo-labeled certain samples is a necessary step in SHOT and involves backpropagation for $N_{\text{itr}}$ iterations. Since each backpropagation iteration is twice as complex as a forward CNN pass, this model adaptation step has a complexity of $2N_{\text{itr}}$ times that of the forward pass, as mentioned in Table V.
Matched field processing method in (3) involves two complex matrix multiplications. Since $\mathbf{\tilde{r}}(f,d)$ is of size $L\times 1$, $d\in\{d_{1},...,d_{N_{\text{tr}}}\}$, and $\tilde{\mathbf{C}}$ is of size $L\times L$, this yields a memory complexity of $2N_{\text{tr}}L$ floating point numbers and a computational complexity of $N_{\text{tr}}(8L^{2}+8L)$. Table V summarizes the complexities of different modules.
TABLE V: Memory and Computational Complexity Comparison
Model Computational Complexity Memory Complexity
MFP $N_{\text{tr}}(8L^{2}+8L)$ $2N_{\text{tr}}L$
CNN-c forward pass $(87,616+80N_{\phi})L^{2}+2MN_{\phi}$ $(86+40N_{\phi})L^{2}+MN_{\phi}+43,808$
JSEA-c refinement $O(M+MW+N_{P}|\mathcal{S}|)$ $N_{\text{test}}(L+1)$
JSEA-c or SHOT model adaptation $2N_{\text{itr}}((87,616+80N_{\phi})L^{2}+2N_{c}N_{\phi})$ $N_{\text{test}}L+(86+40N_{\phi})L^{2}+MN_{\phi}+43,808$